# reference.bib

### v1

@@ 5.1 Experimental Setup
Multi-Modal Datasets. We collect a multi-modal dataset from real-world financial information to construct a representation of the market environment. It is composed of stock price information, daily financial news, company filing reports, and ECC audio recordings ranging from August 1, 2020, to August 14, 2023, as detailed in Appendix. A.1. Each data source features distinct timeliness and is assigned to separate analyst agents to process. Typically, the time-sensitivity of annual filing reports (Form 10K’s) is more persistent, that of quarterly filing reports (Form 10Q’s) and ECC are secondary, and the daily financial news is the most instantaneous.
Evaluation Metrics. We evaluate the performance of FinCon and compare it with other state-of-the-art (SOTA) LLM-based and DRL-based agent systems using three standard financial performance metrics: Cumulative Return (CR%) hull2007risk, the Sharpe Ratio (SR) sharpe1994sharpe and Max Drawdown (MDD%) ang2003downside. Detailed definitions can be found in Appendix A.3.
Comparative Methods. For the single stock trading task, we assess FinCon’s performance against seven algorithmic agents and the Buy-and-Hold (B & H) strategy, a widely accepted baseline. These include three DRL-based agents from the FinRL framework liu2020finrl: A2C, PPO, and DQN, and four SOTA LLM-based agents: Generative Agent by Park et al. park2023generative, FinGPT liu2023fingpt, FinMem yu2023finmem, and FinAgent zhang2024finagent. For the portfolio management task, FinCon competes against the classical Markowitz MV portfolio selection strategy markowitz1952portfolio, the RL-based FinRL-A2C agent liu2020finrl and the B & H strategy that is to hold an equal-weighted position in all trading assets (so-called equal-weighted ETF). Our focus on classical portfolio strategy, RL methods and B & H methods is due to the limited availability of mature LLM agents for portfolio management. The detailed experiment parameter configuration is detailed articulated in Appendix. A.7.
@@ 5.2 Main Results
5.2.1 Single Asset Trading Task
@@
5.2.2 Portfolio Management Task
To address the first research question, we compare FinCon’s performance with the Markowitz Mean-Variance (MV) portfolio markowitz2000mean and FinRL liu2020finrl in managing a small portfolio of TSLA, MSFT, and PFE. Those assets are chosen by the stock selection agent from 42 stocks with sufficient amount of news data (over $800$ new messages in the training and test periods together),referring to Figure 9 in Appendix A.2. The training & testing periods and backbone algorithm are consistent with those used in the single-asset trading task. For the Markowitz MV portfolio, we estimate the covariance matrix and expected returns using the same training data. For FinRL, we use 5-year training data before the test period. Our results, detailed in Table 3, show that FinCon outperforms both the Markowitz MV portfolio and FinRL, achieving significantly higher cumulative returns and Sharpe Ratio, and lower Maximum Drawdown (MDD).
@@ 5.3.2 The Efficacy of Over-Episode Belief Updates Using CVRF
Updated Investment Belief Contents First Update:
{’historical momentum’: ’Enhance the use of momentum indicators by establishing systematic rules for entry and exit based on momentum values to improve consistency and predictability in trading actions.’,
’news insights’: ’Integrate advanced real-time news sentiment analysis to better understand immediate market reactions and adjust trading strategies accordingly…’,
’Form 10-Q’: ’Incorporating annual report insights (10-K) could provide a comprehensive view of the company’s long-term performance and strategic direction, aiding in more informed decision-making.’…,
’other aspects’: [’sector trends’, ’technological advancements’], …}
Last Update:
@@
’news insights’: ’It is recommended to further develop a systematic approach to quantify the impact of news, especially the sentiment scores and regulatory changes, to refine trading decisions.’,
’Form 10-Q’: ’It is suggested that even if there is a strong financial performance present 10-Q reports, it is better to prioritize immediate market signals. For future strategies, it is suggested to balance these aspects more evenly, especially in stable or bullish market conditions, to avoid premature exits from potentially profitable positions.’,…,
’other aspects’: [’sector trends’, ’technological advancements’, ’macroeconomic factors’, ’regulatory risks’], …}
Tradng Action Overlapping Percentages 1. Second vs. First Training Episode:   $46.939\%$
2. Third vs. Second Training Episode: $71.429\%$
3. Fourth vs. Third Training Episode:  $81.633\%$
@@ 6 Conclusion
[…]
Appendix A Appendix
A.1 Raw Data Sources
We assessed the performance of FinCon using multi-modal financial data from August 15, 2020, to August 15, 2023, sourced from reputable databases and APIs including Yahoo Finance (via yfinance), Alpaca News API, and Capital IQ, detailed explained in Table. These data, initially stored in the Raw Financial Data Warehouse as available observations of the financial market environment, are diverged into the corresponding FinCon’s Analysts’ Procedural Memory Databases based on timeliness through working memory’s summarization operation.
Data Sources News data associated with ticker: News data is sourced from REFINITIV REAL-TIME NEWS mainly contains news from Reuters. Form 10-Q, Part 1 Item 2 (Management’s Discussion and Analysis of Financial Condition and Results of Operations): Quarterly reports (Form 10-Q) are required by the U.S. Securities and Exchange Commission (SEC). Form 10-k, Section 7 (Management’s Discussion and Analysis of Financial Condition and Results of Operations): Annual reports (Form 10-K) are required by the U.S. Securities and Exchange Commission (SEC), sourced from EDGAR, and downloaded via SEC API. Historical stock price: Daily open price, high price, close price, adjusted close price, and volume data from Yahoo Finance. Zacks Equity Research: Zacks Rank: The Zacks Rank is a short-term rating system that is most effective over the one- to three-month holding horizon. The underlying driver for the quantitatively-determined Zacks Rank is the same as the Zacks Recommendation and reflects trends in earnings estimate revisions. Zacks Analyst: Reason to Sell, Reason to Buy, and potential risks. Earning Conference Calls (ECC): ECC is a type of unstructured financial data (audio) that is crucial for understanding market dynamics and investor sentiment. The company executive board delivers ECC about recent financial outcomes, future projections, and strategic directions. Recent studies have underscored the importance of not only the textual content of these calls but also the audio feature. Analyses have revealed that the audio elements—such as tone, pace, and inflections—offer significant predictive value regarding company performance and stock movements qin2019you; yang2020html.
Table 6: Raw data and memory warehouses of FinCon
A.2 Distribution of Data
Figure 9: The distribution of news from REFINITIV REAL-TIME NEWS for the 42 stocks in the experiments
Figure 10: The distribution of 10k10q from Securities and Exchange Commission (SEC) for the 42 stocks in the experiments
Figure 11: The distribution of Analyst Report from Zacks Equity Research for the 42 stocks in the experiments
A.3 Formula of Classic Financial Metrics for Risk Estimator
Profit and Loss (PnL)hull2017options: PnL quantifies the net outcome of trading activities over a specified period by accounting for the realized gains and losses from financial instruments like stocks and derivatives.
Value at Risk (VaR) of PnLhull2017options: VaR is a statistical tool used to estimate the potential loss in a portfolio, within a defined confidence interval. Mathematically, it is defined as Equation 2:
$\text{VaR}_{\alpha}(PnL)=\inf\left\{l\in\mathbb{R}:\mathbb{P}(PnL\leq l)\geq\alpha\right\}$ (2)
where $\alpha$ is the confidence level.
Conditional Value at Risk (CVaR) of PnLhull2017options: CVaR is a statistical tool used to estimate the expected potential loss worse than the VaR value in a portfolio, within a defined confidence interval. Mathematically, it is defined as Equation 3:
$\text{CVaR}_{\alpha}(PnL)=\mathbb{E}\Big{\{}PnL|PnL\leq\text{VaR}_{\alpha}(PnL)\Big{\}}$ (3)
where $\alpha$ is the confidence level.
Cumulative Return of PnL hull2007risk: Cumulative Return is a key trading performance metric because it provides a comprehensive insight into investment performance, especially for strategies that emphasize long-term growth and reinvestment. The effectiveness of different investment strategies is evaluated based on their Cumulative Returns, which reflect the total change in value over time. In this study, we compute Cumulative Returns over the specified period by summing daily logarithmic returns, as outlined in Equation 4. This method is widely accepted in the finance area due to its ability to precisely capture minor price fluctuations and symmetrically address gains and losses. In essence, a higher Cumulative Return typically indicates a more effective strategy.
Cumulative Return $=\sum_{t=1}^{n}r_{i}$
$=\sum_{t=1}^{n}\left[\ln\left(\frac{p_{t+1}}{p_{t}}\right)\cdot\text{action}_{t}\right],$ (4)
where $r_{i}$ represents the PnL for day $t+1$, $p_{t}$ is the closing price on day $t$, $p_{t+1}$ is the closing price on day $t+1$, and $\text{action}_{t}$ denotes the trading decision made by the model for that day.
Sharpe Ratio of PnL sharpe1994sharpe: Sharpe Ratio is another core metric for evaluating investment performance and adjusting returns for risk. It is calculated by dividing the portfolio’s average PnL ($R_{p}$) over the risk-free rate ($R_{f}$) by its volatility ($\sigma_{p}$), as shown in Equation 5. This metric adjusts returns for risk, with a higher ratio indicating better risk-adjusted performance. Essential in comparing different portfolios or strategies, it contextualizes performance against similar investments. Although a Sharpe Ratio above 1 is typically considered favorable and above 2 as excellent, these benchmarks can vary depending on the context of comparison.
Sharpe $\textbf{Sharpe Ratio}=\frac{R_{p}-R_{f}}{\sigma_{p}}$ (5)
Max Drawdown of PnL ang2003downside: Max Drawdown is a metric for assessing risk. It represents the most significant decrease in a portfolio’s value, from its highest ($P_{\text{peak}}$) to its lowest point ($P_{\text{trough}}$) until a new peak emerges, detailed in Equation 6. Indicative of investment strategy robustness, a smaller Max Drawdown suggests reduced risk.
Max $\textbf{Max Drawdown}=\text{max}(\frac{P_{\text{peak}}-P_{\text{trough}}}{P_{\text{peak}}})$ (6)
A.4 Baseline and Comparative Models on Single Stock Trading Task
Buy-and-Hold strategy (B&H):
A passive investment approach, where an investor purchases stocks and holds onto them for an extended period regardless of market fluctuations, is commonly used as a baseline for comparison of stock trading strategies.
DRL trading agents:
As the FinMem is practiced and examined on the basis of single stock trading and discrete trading actions, we choose three advanced DRL algorithms fitting into the same scenarios according to the previous and shown expressive performance in the work of Liu et al liu2021finrl; liu2022finrl. The DRL training agents only take numeric features as inputs.
  • Proximal Policy Optimization (PPO): PPO (schulman2017proximal) is employed in stock trading due to its stability and efficiency. One salient advantage of PPO is that it maintains a balance between exploration and exploitation by bounding the policy update, preventing drastic policy changes.
  • Deep Q-Network (DQN): DQN (mnih2013playing) is an adaptation of Q-learning, that can be used to optimize investment strategies. Unlike traditional Q-learning that relies on a tabular approach for storing Q-values, DQN generalizes Q-value estimation across states using deep learning, making it more scalable for complex trading environments.
LLM trading agents:
We evaluate FinCon against four LLM agents in the context of stock trading.
  • General-purpose Generative Agents – GA:The generative AI agent by Park et al. park2023generative, originally intended to simulate realistic human behavior and make everyday decisions, has been adapted here for specific stock trading tasks. This agent’s architecture includes a memory module that employs recency, relevance, and importance metrics to extract pivotal memory events for informed decision-making. However, it does not provide a layered memory module to effectively differentiate the time sensitivities unique to various types of financial data. Additionally, although it features a profiling module to define agent attributes like professional background, the model does not specify the agent’s persona. In our experiments, we modified the original prompt template created by Park et al., which was intended for general daily tasks, to suit financial investment tasks.
  • FinGPT: A novel open-source LLM framework specialized for converting incoming textual and numeric information into informed financial decision-making, introduced by Yang et alyang2023fingpt. It claims superiority over the traditional buy-and-hold strategy.
  • FinMem:FinMem employs a specialized profiling module and self-adaptive risk settings for enhanced market robustness. Its memory module integrates working memory and layered long-term memory, enabling effective data processing. This allows FinMem to leverage market insights and improve trading decisions yu2023finmem.
  • FinAgent:FinAgent developed upon FinMem, which leverages the use of tool-using capabilities of LLMs to incorporate multi-modal financial data zhang2024finagent. It claims an further improved trading performance on single asset trading (stocks and cryptocurrencies).
A.5 Portfolio Management
Markowitz Portfolio Selection: introduced by Harry Markowitz in 1952 markowitz1952portfolio, is a framework for constructing portfolios that optimize expected return for a given level of risk or minimize risk for a given level of expected return. This method uses expected returns, variances, and covariances of asset returns to determine the optimal asset allocation, thereby balancing risk and return through diversification.
FinRL-A2C: is an RL algorithm proposed to address single stock trading and portfolio optimization problems in liu2020finrl. The RL models makes trading decisions (i.e., portfolio weights) based on the observation of previous market conditions and the brokerage information of the RL-agent. The implementation of this algorithm  [^1] https://github.com/AI4Finance-Foundation/FinRL-Meta is provided and is used as baselines in our study.
A.6 Ranking Metrics for Procedural Memory in FinCon
Upon receiving an investment inquiry, each agent in FinCon retrieves the top-$K$ pivotal memory events from its procedural memory, where $K$ is a hyperparameter. These events are selected based on their information retrieval score. For any given memory event $E$, its information retrieval score $\gamma^{E}$ is defined by
$\gamma^{E}=S_{\text{Relevancy}}^{E}+S_{\text{Importance}}^{E}$ (7)
which is adpated from Park et al park2023generative but with modified relevancy and importance computations, and is scaled to $[0,1]$ before summing up. Upon the arrival of a trade inquiry $P$ in processing memory event $E$ via LLM prompts, the agent computes the relevancy score $S_{\text{Relevancy}}^{E}$ that measures the cosine similarity between the embedding vectors of the memory event textual content $\mathbf{m_{E}}$ and the LLM prompt query $\mathbf{m_{P}}$, which is defined as follows:
$S_{\text{Relevancy}}^{E}=\frac{\mathbf{m_{E}}\cdot\mathbf{m_{P}}}{\|\mathbf{m_{E}}\|_{2}\times\|\mathbf{m_{P}}\|_{2}}$ (8)
Note that the LLM prompt query inputs trading inquiry and trader characteristics. On the other hand, the importance score $S_{\text{Importance}}^{E}$ is inversely correlates with the time gap between the inquiry and the event’s memory timestamp $\delta t=t_{\text{P}}-t_{E}$, mirroring Ebbinghaus’s forgetting curve murre2015replication. More precisely, if we denote the initial score value of memory event $v^{E}$ and degrading ratio $\theta\in(0,1)$, then the importance score is computed via
$S_{\text{Importance}}^{E}=v^{E}\times\theta^{\delta t}$ (9)
Note that the ratio $\theta$ measures the diminishing importance of an event over time, which is inspired by design of park2023generative. But in our design, the factors of recency and importance are handled by one equation. Different agents in FinCon admit different choices of $\{v^{E},\theta\}$ for memory event $E$.
Additionally, an access counter function facilitates memory event augmentation, so that critical events impacting trading decisions can be augmented by FinCon, while trivial events are gradually faded. This is achieved by using the LLM validation tool Guardrails AI guardrailsai to track critical memory ID. A memory ID deemed critical to investment gains receives $+5$ to its importance score $S_{\text{Importance}}^{E}$. This access counter implementation enables FinCon to capture and prioritize crucial events based on type and retrieval frequency.
A.7 Detailed Configurations in Experiments
The training period was chosen to account for the seasonal nature of corporate financial reporting and the duration of data retention in FinCon’s memory module. The selected training duration ensures the inclusion of at least one publication cycle of either Form 10-Q, ECC, or Form 10-K. This strategy ensures that the learned conceptualized investment guidance considers a more comprehensive scope of factors. Additionally, the training duration allowed FinCon sufficient time to establish inferential links between financial news, market indicators, and stock market trends, thereby accumulating substantial experience. Furthermore, we set the number of top memory events retrieved for each agent at 5. We ran FinCon. The reported performance outcomes are based on the setting that achieved the highest cumulative return during the testing phase.
To maintain consistency in the comparison, the training and testing phases for the other two LLM-based agents were aligned with those of FinMem. For parameters of other LLM-based agents that are not encompassed by FinMem’s configuration, they were kept in accordance with their original settings as specified in their respective source codes.
FinMem’s performance was benchmarked against that of the most effective comparative model, using Cumulative Return and Sharpe Ratio as the primary evaluation metrics. The statistical significance of FinMem’s superior performance was ascertained through the non-parametric Wilcoxon signed-rank test, which is particularly apt for the non-Gaussian distributed data.

### v2

@@ 5.1 Experimental Setup
Multi-Modal Datasets. We collect a multi-modal dataset from real-world financial information to construct a representation of the market environment. It is composed of stock price information, daily financial news, company filing reports, and ECC audio recordings ranging from August 1, 2020, to August 14, 2023, as detailed in Appendix. 7.1. Each data source features distinct timeliness and is assigned to separate analyst agents to process. Typically, the time-sensitivity of annual filing reports (Form 10K’s) is more persistent, that of quarterly filing reports (Form 10Q’s) and ECC are secondary, and the daily financial news is the most instantaneous.
Evaluation Metrics. We evaluate the performance of FinCon and compare it with other state-of-the-art (SOTA) LLM-based and DRL-based agent systems using three standard financial performance metrics: Cumulative Return (CR%) [73], the Sharpe Ratio (SR) [74] and Max Drawdown (MDD%) [75]. Detailed definitions can be found in Appendix 7.3.
Comparative Methods. For the single stock trading task, we assess FinCon’s performance against seven algorithmic agents and the Buy-and-Hold (B & H) strategy, a widely accepted baseline. These include three DRL-based agents from the FinRL framework [76]: A2C, PPO, and DQN, and four SOTA LLM-based agents: Generative Agent by Park et al. [11], FinGPT [77], FinMem [23], and FinAgent [24]. For the portfolio management task, FinCon competes against the classical Markowitz MV portfolio selection strategy [1], the RL-based FinRL-A2C agent [76] and the B & H strategy that is to hold an equal-weighted position in all trading assets (so-called equal-weighted ETF). Our focus on classical portfolio strategy, RL methods and B & H methods is due to the limited availability of mature LLM agents for portfolio management. The detailed experiment parameter configuration is detailed articulated in Appendix. 7.7.
@@ 5.2.1 Single Asset Trading Task
5.2.2 Portfolio Management Task
To address the first research question, we compare FinCon’s performance with the Markowitz Mean-Variance (MV) portfolio [78] and FinRL [76] in managing a small portfolio of TSLA, MSFT, and PFE. Those assets are chosen by the stock selection agent from 42 stocks with sufficient amount of news data (over $800$ new messages in the training and test periods together), referring to Figure 9 in Appendix 7.2. The training & testing periods and backbone algorithm are consistent with those used in the single-asset trading task. For the Markowitz MV portfolio, we estimate the covariance matrix and expected returns using the same training data. For FinRL, we use 5-year training data before the test period. Our results, detailed in Table 3, show that FinCon outperforms both the Markowitz MV portfolio and FinRL, achieving significantly higher cumulative returns and Sharpe Ratio, and lower Maximum Drawdown (MDD).
@@ 5.3.2 The Efficacy of Over-Episode Belief Updates Using CVRF
Updated Investment Belief Contents First Update: {’historical momentum’: ’Enhance the use of momentum indicators by establishing systematic rules for entry and exit based on momentum values to improve consistency and predictability in trading actions.’, ’news insights’: ’Integrate advanced real-time news sentiment analysis to better understand immediate market reactions and adjust trading strategies accordingly…’, ’Form 10-Q’: ’Incorporating annual report insights (10-K) could provide a comprehensive view of the company’s long-term performance and strategic direction, aiding in more informed decision-making.’…, ’other aspects’: [’sector trends’, ’technological advancements’], …} Last Update:
@@
’news insights’: ’It is recommended to further develop a systematic approach to quantify the impact of news, especially the sentiment scores and regulatory changes, to refine trading decisions.’, ’Form 10-Q’: ’It is suggested that even if there is a strong financial performance present 10-Q reports, it is better to prioritize immediate market signals. For future strategies, it is suggested to balance these aspects more evenly, especially in stable or bullish market conditions, to avoid premature exits from potentially profitable positions.’,…, ’other aspects’: [’sector trends’, ’technological advancements’, ’macroeconomic factors’, ’regulatory risks’], …}
Tradng Action Overlapping Percentages 1. Second vs. First Training Episode:   $46.939\%$ 2. Third vs. Second Training Episode: $71.429\%$ 3. Fourth vs. Third Training Episode:  $81.633\%$